# Kinematic cosmic dipole from a large sample of strong lenses

### 11 Mar 2026 v1

@@ 6 Conclusion
  • The inclusion of kinematic information, either through spectroscopic velocity-dispersion measurements or via FP relations, significantly improves the constraining power of the method. Even in the absence of spectroscopic redshift information, the combination of FP relations and photometric redshifts leads to an improvement of $\sim 50$% compared to lensing-only constraints.
  • Already existing or planned spectroscopic follow-up observations from 4MOST [4SLSLS] and DESI [e.g., Huang2025] should yield a discriminating power at the level of $\sim 2\sigma$ between the value of $v_{o}$ measured from the CMB and that measured from sources number counts.
  • A hypothetical dedicated strong-lensing spectroscopic follow-up of the full Euclid sample could provide a decisive contribution to this debate, with a discriminating power reaching $\sim 4\sigma$.
@@ Acknowledgments
MM acknowledges support from the Swiss National Science Foundation (SNSF) through return CH grant P5R5PT_225598 and Ambizione grant PZ00P2_223738. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (LensEra: grant agreement No 945536). TEC is funded by the Royal Society through a University Research Fellowship. CB acknowledges support from the Swiss National Science Foundation (SNSF).
@@ Appendix A Calculation of the mean Einstein radius integrated over redshifts
The component of $\theta^{\prime}_{E}$ along the $\bm{e}^{\prime}_{1}$ direction at fixed $z^{\prime}_{s}$ and $z^{\prime}_{l}$ is given by
$\theta^{\prime}_{E1}(\delta^{\prime},\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=\theta^{\prime}_{E1}(\delta^{\prime},\bm{\hat{n}}^{\prime},r^{\prime}(z^{\prime}_{s},\bm{\hat{n}}^{\prime}),r^{\prime}(z^{\prime}_{l},\bm{\hat{n}}^{\prime}))\,$, (A.3)
@@
Doing the same split for the redshift of the lens, we can insert them into (A.3) and Taylor expand at linear order in $v_{o}/c$
$\theta^{\prime}_{E1}(\delta^{\prime},\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=\theta_{E}(\delta^{\prime},r(z^{\prime}_{s}),r(z^{\prime}_{l}))\cos\delta^{\prime}+\left(\frac{\partial\theta_{E}}{\partial r_{s}{\cal H}(z^{\prime}_{s})}+\frac{\partial\theta_{E}}{\partial r_{l}{\cal H}(z^{\prime}_{l})}\right)\cos\delta^{\prime}\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\,.$ (A.6)
From this we see that even though the angles along $\bm{e}^{\prime}_{1}$ are not aberrated, the fact that the redshifts of the lens and of the source are affected by the boost has an impact on $\theta^{\prime}_{E1}$. Physically, this is due to the fact that a source at fixed redshift $z^{\prime}_{s}$ is not at the same physical distance from the observer if it is in the direction of the boost or if it is in the opposite direction. Since the Einstein radius depends on the distance of the source, this impacts the observed $\theta^{\prime}_{E1}$. The calculation for $\theta^{\prime}_{E2}$ is similar, with the effect from the redshift added to the effect of aberration
$\theta^{\prime}_{E2}(\delta^{\prime},\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=$ $\theta_{E}(\delta^{\prime},r(z^{\prime}_{s}),r(z^{\prime}_{l}))\sin\delta^{\prime}\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$ (A.7)
$+\left(\frac{\partial\theta_{E}}{\partial r_{s}{\cal H}(z^{\prime}_{s})}+\frac{\partial\theta_{E}}{\partial r_{l}{\cal H}(z^{\prime}_{l})}\right)\sin\delta^{\prime}\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\,.$
Computing $\theta^{\prime}_{E}(\delta^{\prime},\bm{\hat{n}}^{\prime},z_{s}^{\prime},z_{l}^{\prime})$ and averaging over $\delta^{\prime}$, we find
$\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=$ $\theta_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)+\left((1+z_{z})\frac{\partial\theta_{E}}{\partial z_{s}}+(1+z_{l})\frac{\partial\theta_{E}}{\partial z_{l}}\right)\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\,$, (A.8)
where we have used that $\partial/\partial r={\cal H}(1+z)\partial/\partial z$.
@@
$=\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}))}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}}\theta_{E}\left(1+2\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
The terms in the last line can directly be integrated, giving rise to boundary terms that exactly vanish since the number of rings at redshift 0 and at infinity is zero. The only effect that survives is therefore the effect of aberration, which modifies both the number of rings (second term in the second line of (A.11)) and the Einstein radius (third term in the second line of (A.11)). A similar calculation can be performed for the denominator of Eq. (2.9), $dN(\bm{\hat{n}}^{\prime})/d\Omega^{\prime}$, leading to
@@
Taking the ratio of Eqs. (A.11) and (A.12) , we obtain for the mean
$\langle\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime})\rangle=\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}f(z^{\prime}_{s},z^{\prime}_{l})\theta_{E}(z^{\prime}_{s},z^{\prime}_{l})\,$, (A.13)
@@ Appendix B Impact of the selection function on the dipole
We compute here the impact of selection in flux and size on the dipole. We work in redshift space, but as we will see, we would obtain the same result using comoving distances, once the signal is integrated from 0 to $\infty$.
@@
$\theta_{E}^{\prime}=\theta^{\rm int}_{E}\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)g(z_{s}^{\prime},z^{\prime}_{l})\,$, (B.4)
@@
$\theta^{\rm int}_{E*}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))+\delta\theta^{\rm int}_{E*}(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=\frac{\theta^{\prime}_{E*}}{g(r(z_{s}^{\prime}),r(z^{\prime}_{l}))}\left(1+\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}-\frac{\partial g}{\partial r_{s}}\frac{\delta r_{s}}{g}-\frac{\partial g}{\partial r_{l}}\frac{\delta r_{l}}{g}\right)\,.$ (B.5)
@@
$=\int_{\theta^{\rm int}_{E*}}\mathrm{d}\theta^{\rm int}_{E}\theta^{\rm int}_{E}\Bigg\{\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}),\theta^{\rm int}_{E},L>L_{*})}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}\mathrm{d}\theta^{\rm int}_{E}}g(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1+\frac{3}{2}\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
$=\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}\Bigg\{\int_{\theta^{\rm int}_{E*}}\mathrm{d}\theta^{\rm int}_{E}\theta^{\rm int}_{E}\Bigg[\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}),\theta^{\rm int}_{E},L>L_{*})}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}\mathrm{d}\theta^{\rm int}_{E}}g(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1+\frac{3}{2}\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
$\delta\theta^{\rm int}_{E*}=\theta^{\rm int}_{E*}\left(\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}-\frac{\partial g}{\partial r_{s}}\frac{\delta r_{s}}{g}-\frac{\partial g}{\partial r_{l}}\frac{\delta r_{l}}{g}\right)\,$, (B.11)
@@
$=\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}\int_{\theta^{\rm int}_{E*}}\mathrm{d}\theta^{\rm int}_{E}\theta^{\rm int}_{E}\Bigg[\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}),\theta^{\rm int}_{E},L>L_{*})}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}\mathrm{d}\theta^{\rm int}_{E}}g(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1+\frac{3}{2}\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
$\langle\theta^{\prime}_{E}\rangle$ $=\langle\theta_{E}\rangle\left(1-\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}\right)$ (B.15)
@@
$+\frac{v_{o}}{2c}\cos\theta^{\prime}_{\rm cm}(\langle\theta_{E}\rangle-\theta_{E*})\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}f_{2}(z^{\prime}_{s},z^{\prime}_{l},\theta_{E}>\theta_{E*},L>L_{*})\mu(z^{\prime}_{s},z^{\prime}_{l},\theta_{E*},L_{*})\,.$
@@
These two distributions are unboosted quantities, which depend on the unboosted distance $r$, the unboosted angle $\theta_{E}$, and the intrinsic unboosted luminosity threshold $L_{*}$. In practice, however, to predict the amplitude of the second and third lines in Eq. (B.15), we can directly use the measured mean redshift distribution as a function of the observed redshifts, the observed angle $\theta^{\prime}_{E}$, and the observed flux threshold $F^{\prime}>F^{\prime}_{*}$, since $f_{1}$ and $f_{2}$ multiply already $v_{o}/c$.
@@ Appendix D Observer velocity effect on the Fundamental Plane observables
$\alpha=\alpha^{\prime}\left(1+\frac{v_{o}}{2c}\cos(\theta_{\rm cm}^{\prime})\right)\,.$ (D.1)
@@
$R_{e}=\alpha d_{\rm A}=\alpha^{\prime}d[z_{l}^{\prime}]\left(1+\frac{v_{o}}{2c}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{d[z_{l}^{\prime}]}\frac{v_{o}}{c}\right)\,.$ (D.2)
@@
$F=\int_{\mathbb{R}_{+}}\mathrm{d}\nu\mathrm{d}^{2}\int_{\mathcal{S}}\bm{\hat{n}}\,I(\nu,\bm{\hat{n}})R(\nu)\,.$ (D.4)
@@
$=\log(\alpha^{\prime}d[z_{l}^{\prime}])-b\log(I_{e}^{\prime})-c+\log\left(1+\frac{1}{2}\bm{v}_{o}\cdot\bm{\hat{n}^{\prime}}+\frac{D_{L}}{d[z_{l}^{\prime}]}\frac{v_{o}}{c}\right)\,.$ (D.13)
@@
$\Sigma_{v}$ $=\left(\frac{1}{2a}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{ad[z_{l}^{\prime}]}\right)\,.$ (D.16)
@@
$\theta^{\prime}_{E,\rm FP}=\frac{4\pi\sigma_{v}^{\prime 2}}{c^{2}}\frac{d[z_{l}^{\prime},z_{s}^{\prime}]}{d[z_{s}^{\prime}]}\left(1-\frac{1}{2}\frac{v_{o}}{c}\cos(\theta^{\prime}_{\rm cm})+\frac{D_{LS}}{d[z_{l}^{\prime},z_{s}^{\prime}]}\frac{v_{o}}{c}-\frac{D_{S}}{d[z_{s}^{\prime}]}\frac{v_{o}}{c}+2\Sigma_{v}\frac{v_{o}}{c}\right)\,.$ (D.18)

### 19 Mar 2026 v2

@@ 6 Conclusion
  • The inclusion of kinematic information, either through spectroscopic velocity-dispersion measurements or via FP relations, significantly improves the constraining power of the method. Even in the absence of spectroscopic redshift information, the combination of FP relations and photometric redshifts leads to an improvement of $\sim 40$% compared to lensing-only constraints.
  • Already existing or planned spectroscopic follow-up observations from 4MOST [4SLSLS] and DESI [e.g., Huang2025] should yield a discriminating power at the level of $\sim 2.8\sigma$ between the value of $v_{o}$ measured from the CMB and that measured from source number counts.
  • A hypothetical dedicated strong-lensing spectroscopic follow-up of the full Euclid sample could provide a decisive contribution to this debate, with a discriminating power reaching $\sim 5\sigma$.
@@ Acknowledgments
We are grateful to Koen Kuijken for identifying a calculation error in an earlier version of this manuscript. MM acknowledges support from the Swiss National Science Foundation (SNSF) through return CH grant P5R5PT_225598 and Ambizione grant PZ00P2_223738. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (LensEra: grant agreement No 945536). TEC is funded by the Royal Society through a University Research Fellowship. CB acknowledges support from the Swiss National Science Foundation (SNSF).
@@ Appendix A Calculation of the mean Einstein radius integrated over redshifts
The Einstein radius $\theta^{\prime}_{E}$ at fixed $z^{\prime}_{s}$ and $z^{\prime}_{l}$ is given by
$\theta^{\prime}_{E}(\delta^{\prime},\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=\theta^{\prime}_{E}(\delta^{\prime},\bm{\hat{n}}^{\prime},r^{\prime}(z^{\prime}_{s},\bm{\hat{n}}^{\prime}),r^{\prime}(z^{\prime}_{l},\bm{\hat{n}}^{\prime}))\,$, (A.3)
@@
Doing the same split for the redshift of the lens, we can insert it into (A.3) and Taylor expand at linear order in $v_{o}/c$
$\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=\theta^{\prime}_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))+\left(\frac{\partial\theta_{E}}{\partial r_{s}{\cal H}(z^{\prime}_{s})}+\frac{\partial\theta_{E}}{\partial r_{l}{\cal H}(z^{\prime}_{l})}\right)\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}$
$=\theta_{E}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)+\left((1+z_{z})\frac{\partial\theta_{E}}{\partial z_{s}}+(1+z_{l})\frac{\partial\theta_{E}}{\partial z_{l}}\right)\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\,$, (A.6)
where in the second line we have used that $\partial/\partial r={\cal H}(1+z)\partial/\partial z$. From this, we see that the fact that the redshifts of the lens and the source are affected by the boost adds another contribution to $\theta^{\prime}_{E}$. Physically, this is due to the fact that a source at a fixed redshift $z^{\prime}_{s}$ is not at the same physical distance from the observer if it is in the direction of the boost or if it is in the opposite direction. Since the Einstein radius depends on the distance of the source, this impacts the observed $\theta^{\prime}_{E}$.
@@
$=\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}))}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}}\theta_{E}\left(1+2\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
The terms in the last line can be directly integrated, giving rise to boundary terms that exactly vanish since the number of rings at redshift 0 and at infinity is zero. The only effect that survives is therefore the effect of aberration, which modifies both the number of rings (second term in the second line of (A.9)) and the Einstein radius (third term in the second line of (A.9)). A similar calculation can be performed for the denominator of Eq. (2.6), $dN(\bm{\hat{n}}^{\prime})/d\Omega^{\prime}$, leading to
@@
Taking the ratio of Eqs. (A.9) and (A.10) , we obtain the mean
$\langle\theta^{\prime}_{E}(\bm{\hat{n}}^{\prime})\rangle=\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}f(z^{\prime}_{s},z^{\prime}_{l})\theta_{E}(z^{\prime}_{s},z^{\prime}_{l})\,$, (A.11)
@@ Appendix B Impact of the selection function on the dipole
We compute here the impact of selection in flux and size on the dipole. We work in redshift space, but as we will see, we would obtain the same result using comoving distances once the signal is integrated from 0 to $\infty$.
@@
$\theta_{E}^{\prime}=\theta^{\rm int}_{E}\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)g(z_{s}^{\prime},z^{\prime}_{l})\,$, (B.4)
@@
$\theta^{\rm int}_{E*}(r(z^{\prime}_{s}),r(z^{\prime}_{l}))+\delta\theta^{\rm int}_{E*}(\bm{\hat{n}}^{\prime},z^{\prime}_{s},z^{\prime}_{l})=\frac{\theta^{\prime}_{E*}}{g(r(z_{s}^{\prime}),r(z^{\prime}_{l}))}\left(1+\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}-\frac{\partial g}{\partial r_{s}}\frac{\delta r_{s}}{g}-\frac{\partial g}{\partial r_{l}}\frac{\delta r_{l}}{g}\right)\,.$ (B.5)
@@
$=\int_{\theta^{\rm int}_{E*}}\mathrm{d}\theta^{\rm int}_{E}\theta^{\rm int}_{E}\Bigg\{\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}),\theta^{\rm int}_{E},L>L_{*})}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}\mathrm{d}\theta^{\rm int}_{E}}g(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1+\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
$=\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}\Bigg\{\int_{\theta^{\rm int}_{E*}}\mathrm{d}\theta^{\rm int}_{E}\theta^{\rm int}_{E}\Bigg[\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}),\theta^{\rm int}_{E},L>L_{*})}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}\mathrm{d}\theta^{\rm int}_{E}}g(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1+\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
$\delta\theta^{\rm int}_{E*}=\theta^{\rm int}_{E*}\left(\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}-\frac{\partial g}{\partial r_{s}}\frac{\delta r_{s}}{g}-\frac{\partial g}{\partial r_{l}}\frac{\delta r_{l}}{g}\right)\,$, (B.11)
@@
$=\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}\int_{\theta^{\rm int}_{E*}}\mathrm{d}\theta^{\rm int}_{E}\theta^{\rm int}_{E}\Bigg[\frac{\mathrm{d}N(r(z^{\prime}_{s}),r(z^{\prime}_{l}),\theta^{\rm int}_{E},L>L_{*})}{\mathrm{d}\Omega\mathrm{d}z_{s}\mathrm{d}z_{l}\mathrm{d}\theta^{\rm int}_{E}}g(r(z^{\prime}_{s}),r(z^{\prime}_{l}))\left(1+\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$
@@
$\langle\theta^{\prime}_{E}\rangle$ $=\langle\theta_{E}\rangle\left(1-\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}\right)$ (B.15)
@@
$+\frac{v_{o}}{c}\cos\theta^{\prime}_{\rm cm}(\langle\theta_{E}\rangle-\theta_{E*})\int\mathrm{d}z^{\prime}_{s}\int\mathrm{d}z^{\prime}_{l}f_{2}(z^{\prime}_{s},z^{\prime}_{l},\theta_{E}>\theta_{E*},L>L_{*})\mu(z^{\prime}_{s},z^{\prime}_{l},\theta_{E*},L_{*})\,.$
@@
These two distributions are unboosted quantities, which depend on the unboosted distance $r$, the unboosted angle $\theta_{E}$, and the intrinsic unboosted luminosity threshold $L_{*}$. In practice, however, to predict the amplitude of the second and third lines in Eq. (B.15), we can directly use the measured mean redshift distribution as a function of the observed redshifts, the observed angle $\theta^{\prime}_{E}$, and the observed flux threshold $F^{\prime}>F^{\prime}_{*}$, since $f_{1}$ and $f_{2}$ already multiply $v_{o}/c$.
@@ Appendix D Observer velocity effect on the Fundamental Plane observables
$\alpha=\alpha^{\prime}\left(1+\frac{v_{o}}{c}\cos(\theta_{\rm cm}^{\prime})\right)\,.$ (D.1)
@@
$R_{e}=\alpha d_{\rm A}=\alpha^{\prime}d[z_{l}^{\prime}]\left(1+\frac{v_{o}}{c}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{d[z_{l}^{\prime}]}\frac{v_{o}}{c}\right)\,.$ (D.2)
@@
$F=\int_{\mathbb{R}_{+}}\mathrm{d}\nu\int_{\mathcal{S}}\mathrm{d}^{2}\bm{\hat{n}}\,I(\nu,\bm{\hat{n}})R(\nu)\,.$ (D.4)
@@
$=\log(\alpha^{\prime}d[z_{l}^{\prime}])-b\log(I_{e}^{\prime})-c+\log\left(1+\bm{v}_{o}\cdot\bm{\hat{n}^{\prime}}+\frac{D_{L}}{d[z_{l}^{\prime}]}\frac{v_{o}}{c}\right)\,.$ (D.13)
@@
$\Sigma_{v}$ $=\left(\frac{1}{a}\cos(\theta_{\rm cm}^{\prime})+\frac{D_{L}}{ad[z_{l}^{\prime}]}\right)\,.$ (D.16)
@@
$\theta^{\prime}_{E,\rm FP}=\frac{4\pi\sigma_{v}^{\prime 2}}{c^{2}}\frac{d[z_{l}^{\prime},z_{s}^{\prime}]}{d[z_{s}^{\prime}]}\left(1-\frac{v_{o}}{c}\cos(\theta^{\prime}_{\rm cm})+\frac{D_{LS}}{d[z_{l}^{\prime},z_{s}^{\prime}]}\frac{v_{o}}{c}-\frac{D_{S}}{d[z_{s}^{\prime}]}\frac{v_{o}}{c}+2\Sigma_{v}\frac{v_{o}}{c}\right)\,.$ (D.18)